OpenStudy (anonymous):

find the sum of all positive multiples of 5 less than 400

OpenStudy (anonymous):

use the concept of Geometric Series example 5,10,15,20,25,...395

OpenStudy (anonymous):

or you can used the standard form for a NON GEOMETRIC SERIES, summ of nth term=a+(n-1)d, where a is first term, d is diference, n is the nth term

OpenStudy (anonymous):

for geometric series you have s = a(1-r^n)/ (1-r)

OpenStudy (anonymous):

just to give you an hint you can use 5+(39-1)*5= summation , figure out why I am suggestion 39

OpenStudy (anonymous):

thanks but i got the answer. 15,800

OpenStudy (anonymous):

And is that the correct answer